1. How many GRAMS of zinc are present in 2.18 grams of zinc chromate ?
1
Expert's answer
2017-10-03T15:01:07-0400
1. m(Zn) = (m(ZnCrO4) * M(Zn)) / M(ZnCrO4) = 2.18 * 65 / 181 = 0.78g